# Relightable Neural Actor with Intrinsic Decomposition and Pose Control

### v1

@@ 5.3 Applications
Figure 8: Our method offers the possibility to edit the material properties and simultaneously perform skeletal pose control, free-viewpoint rendering, and relighting of the edited neural actor.
@@ 6 Conclusion
We conclude that the proposed method for the free-viewpoint rendering of humans enables simultaneous relighting, material editing, and pose control at test time with unprecedented quality and level of detail. Our method demonstrates functionality that no competing method making the same assumptions at training time (i.e. from multi-view RGB video inputs only) can provide. We hope that our method and the proposed dataset will provoke further research in this exciting field, bringing the neural rendering of humans to a new qualitative level and paving the way for next-generation metaverse applications.
Limitations. This work focuses on visual fidelity, and representations faster than NeRF could be tested in our framework in future. Moreover, like several previous techniques for the free-viewpoint rendering of humans and avatar generation [25, 37], the proposed method can currently only deal with tight clothing due to the coarse geometry registration.
Supplementary Material
This document accompanies our paper “Relightable Neural Actor with Intrinsic Decomposition and Pose Control” and includes additional implementation details on the network architectures (Sec. 7) and dataset collection (Sec. 8). The dynamic results of our method are provided in the supplementary video.
7 Network Architectures
7.1 Pose-driven Geometry Model
The neural network architecture of our pose-driven geometry model is implemented based on Neural Actor [25] and is illustrated in Fig. 9. The RGB color output $c$ is activated with the sigmoid function, and the density output $\sigma$ is activated with ReLU. The yellow rectangular blocks represent a single fully connected layer.
Figure 9: Neural network architecture of our pose-driven geometry model. The RGB color branch is only used during the training phase of the geometry model and is discarded during the relighting training and inference. The interpolation of texture features (obtained from the texture map) is performed in the UV coordinate $(u,v)$ given by the projection of 3D points to the human mesh, as introduced in Sec. 3.2. The obtained feature vector $\psi$ (represented in purple) is also used in our UVDeltaNet. The symbol $\sim$”refers to positional encoding as in NeRF [32] and “||||| |” refers to feature-wise concatenation.
7.2 NormalNet Model
The goal of NormalNet in our approach is to perform inpainting and refinement of the normal values sampled from the neural field and aggregated in the UV map. Therefore, we leverage partial convolutions [24] as a building block and design a shallow convolutional neural network (CNN) that is shown in Fig. 10.
Figure 10: Neural network architecture of our NormalNet model. Each rectangular block is a 2D partial convolution. Binary masks are applied to all the feature maps and to the input, as defined in Liu et al. [24]. The symbol $\uparrow$” refers to depth-to-space transformation, where the spatial resolution is increased by a factor of $2$ in each dimension, and “||||| |” refers to feature-wise concatenation.
7.3 VisibilityNet Model
We use a similar approach as in the NormalNet architecture for inpainting and refining the visibility information. However, the visibility is highly dependent on the normal values, since half of the light sources in the environment map are back lit as a function of the normal direction. Therefore, the VisibilityNet model also takes as input the sampled normal maps. The shallow CNN architecture of VisibilityNet is shown in Fig. 11.
Figure 11: Neural network architecture of our VisibilityNet model. Rectangular blocks represent 2D partial convolutions. $\uparrow$” and “||||| |” refer to depth-to-space transformation and feature-wise concatenation, respectively.
Figure 12: Neural network architecture of our UVDeltaNet model.
7.4 UVDeltaNet Model
The architecture of UVDeltaNet, introduced in Sec. 3.3.3, is shown in Fig. 12. This shallow MLP has as objective predict a corrective term in UV space, therefore, it takes as input the local coordinates $(u,v,h)$ and the texture features $\psi$ (see Fig. 9), and predicts a corrective term $(u^{\prime},v^{\prime})$.
8 Dataset Collection
We collected our dataset “Relightable Dynamic Actors” with the goal to provide a real dataset with the same person recorded under different light conditions. For each sequence, the actors were instructed to follow a sequence of $10$ activities defined as: stretch arms, walk in a circle, jogging, stretch legs, talk on the phone, use a tablet, stretch legs up, stretch back, wave hands, freestyle. We use a commercial markerless motion tracking system [41] for obtaining the human motion in 3D from the calibrated multi-view videos. To fit SMPL model [27] to our tracking, we first optimize the pose and shape parameters in the rest pose “T-pose” with a 3D body joint loss between SMPL and the tracked 3D pose based on EasyMocap [1].
We obtained the environment map for each sequence by taking multi-exposure pictures from a mirror sphere. The HDR image of the mirror sphere was obtained with the algorithm from Debevec and Malik [8]. We combined two shots of the mirror sphere from different angles. The combined image was unwrapped to the EXR latitude-longitude format, resulting in the HDR environment map. Two samples from our environment maps are shown in Fig. 13.
Figure 13: Samples from our environment maps obtained from indoor and outdoor sequences. On the left is the light probe obtained with HDR pictures from a mirror sphere and on the right is the processed HDR images converted to a latitude-longitude format (EXR). The environment maps (right) are then resized to $32\times{16}$ pixels to be used in our method.

### v2

@@ 5.3 Applications
Figure 8: Our method allows editing the material properties and simultaneously performing skeletal pose control, free-viewpoint rendering, and relighting of the edited neural actor.
6 Limitations
This work focuses on visual fidelity, and representations faster than NeRF could be tested in our framework in future. Moreover, like several previous techniques for the free-viewpoint rendering of humans and avatar generation [32, 45], the proposed method can currently only deal with tight clothing due to the coarse geometry registration. The coarse mesh registration can also negatively affect the fine details in the face and hand synthesis.
@@ 7 Conclusion
We conclude that the proposed method for the free-viewpoint rendering of humans enables simultaneous relighting, material editing, and pose control at test time while requiring a single multi-view video sequence with static illumination during training. Our method demonstrates functionality that no competing method making the same assumptions at training time (i.e. from multi-view RGB video inputs only) can provide. We hope that our method and the proposed dataset will provoke further research in this exciting field, bringing the neural rendering of humans to a new qualitative level and paving the way for next-generation metaverse applications.
Acknowledgements
This work was funded by the ERC Consolidator Grant 4DRepLy (770784), the Saarbrücken Research Center for Visual Computing, Interaction and AI, and the German Research Foundation (DFG) under Grant No. 468670075.